Sara has a collection of nickels, dimes and quarters worth $30.75. she has 15 more dimes than nickels and 3 times as many quarters as dimes. How many coins of each kind does she have?
Answer by mananth(16949)

Nickels =x
dimes= x+15
Quarters = 3(x+15)
Total = 30.75
conver to cents = 3075
5x+10(x+15)+25*3(x+15)= 3075
5x+10x+150+75x+1125=3075
90x+1275=3075
90x=3075-1275
90x=1800
/90
x= 20 number of nickels
x+15 = 20 +15 = 35 dimes
3(x+15) => 3*35=105 quarters